### v1

# Hybrid Minimum-Seeking in Synergistic Lyapunov Functions: Robust Global Stabilization under Unknown Control Directions

Abstract
We consider the problem of robust global stabilization for a control-affine systems with dynamic uncertainty in the control directions and in the presence of topological obstructions that preclude the existence of smooth global control Lyapunov functions. By leveraging a recent Lie-bracket averaging result for hybrid dynamic inclusions introduced in [1], we propose a novel universal hybrid feedback law that induces robust global practical stability by seeking the minimum point of a collection of suitable synergistic Lyapunov functions. As concrete applications of our results, we present novel hybrid high-frequency high-amplitude feedback laws for the solution of robust global stabilization problems on different types of manifolds under unknown control directions, as well as controllers for obstacle avoidance problems in vehicles characterized by kinematic models describing both holonomic and non-holonomic models. By leveraging Lie-bracket averaging for hybrid systems, we show how the proposed hybrid minimum-seeking feedback laws can overcome lack of controllabiltiy during persistent (bounded) periods of time. Numerical simulation results are presented to illustrate the main results.
keywords:
Hybrid systems, Adaptive systems, Nonlinear Control, Stability and Stabilization.
thanks: This work was supported in part by the grants NSF ECCS CAREER 2305756 and AFOSR YIP: FA9550-22-1-0211. Corresponding Author: Jorge I. Poveda.
1 Introduction
Ensuring robustness in the face of uncertainty is a core challenge in controller design for autonomous systems. This challenge becomes significantly more difficult if the uncertainty lies in the control direction. For instance, a controller operating on incorrect assumptions or information on the control direction can lead to damages in the system, as exemplified by a cyber-physical system experiencing severe performance deterioration if the control gain sign unexpectedly changes. Such a scenario can arise without malicious interference, such as an internal software failure, or from external spoofing attacks by adversarial agents [2, 3, 4]. Moreover, since the control direction might be time-varying and occasionally zero (no controllability) for bounded periods, traditional identification or parameter estimation techniques can be insufficient to address the issue due to lack of excitation in the system’s trajectories. Therefore, designing stabilizing feedback laws that can handle uncertainty in control directions, as well as persistent lack of controllability, is crucial for ensuring the resilience of autonomously operating systems in complex and dynamic environments.
Referred to as the problem of stabilization with an unknown sign of the high-frequency gain, or stabilization under unknown control directions, designing stabilizing feedback laws in the presence of uncertainty on control directions has a long history in the literature of robust adaptive control that dates back to the 1970s [5]. The first solution was proposed by Nussbaum [6], which became the common approach to solving this problem. However, it is well-known, e.g. see [7], that the Nussbaum approach can lead to poor performance if the control sign is dynamic. An alternative model-free approach for solving this difficult problem was introduced by Scheinker and Krstic in [7, 8]. The approach introduced in [7] is based on seeking the minimum of a suitably constructed control Lyapunov-like function $V$, using high-frequency oscillatory feedback that relies only on evaluations of this function. As shown in [7, 8], this model-free approach is impervious to persistent changes in the sign of the control gain, which makes it an ideal candidate for the design of stabilizing feedback laws operating in uncertain and adversarial environments. Since its conception in [7], the minimum-seeking approach has found numerous applications, including in the context of 2-D Vehicle Control [8], output regulation problems in nonlinear systems [9], etc, see [10, Sec. 6] for a recent survey.
A core theoretical development that enabled the results of [7, 8] was the emergence of Lie-bracket averaging for ODEs as a powerful framework for the analysis and design of model-free control and optimization algorithms for control-affine systems [11, 12, 13, 14]. Nevertheless, given that the feedback law studied in [7, 8] is continuous, and its stability properties are inherited from the stability properties of its Lie-bracket average system -which is also continuous- the controller inherits the traditional limitations that apply continuous feedback laws, see [15]. Namely, when the system operates in topological spaces that are not contractible to a point, as in, e.g., smooth compact manifolds, the control law cannot achieve robust global stabilization of a point. Such type of topological obstructions to robust global stabilization are also relevant for other applications, such as the problem of obstacle avoidance [16, 17], rigid body global attitude stabilization [18, 19], stabilization on $\mathbb{S}^{1}$, etc. On the other hand, when the sign in the control direction is known, such stabilization problems can be tackled via hybrid control using the so-called synergistic hybrid feedback approach [20, 17, 21]. However, existing results on synergistic hybrid control cannot cope with uncertainty in the control directions. Indeed, even the state-of-the-art results on robust synergistic hybrid feedback for uncertain systems [22] are inapplicable when the control gain sign is uncertain.
The main contribution of this paper is to provide a solution for robust global stabilization under unknown control directions despite topological obstructions to global stability. Specifically, we propose a class of hybrid and oscillatory feedback laws that leverage a collection of suitable (local) control Lyapunov functions. These functions work together to control the system when a global cLf is unavailable. Building on existing results in model-based hybrid control [20, 19, 17] and using a Lie-bracket averaging result for hybrid inclusions established in [1], we establish semi-global practical asymptotic stability for systems with unbounded operational sets and global practical asymptotic stability for systems on smooth compact manifolds, even with unknown control directions that may vanish intermittently. Subsequently, we apply our results to various control problems with unknown control directions where a robustly globally stabilizing continuous control law does not exist even if the control direction is known, including stabilization problems on $\mathbb{S}^{1}$, $\mathbb{S}^{2}$, $SO(3)$, and stabilization with obstacle avoidance for holonomic and non-holonomic vehicles. In all applications, we show how our results open the door for the design of “model-free” controllers based on existing well-posed model-based hybrid algorithms [23, 21, 17, 24], thus demonstrating how the synergistic use of hybrid control [21] and Lie-bracket averaging [1] can simultaneously overcome uncertainties in control directions and the topological obstructions to global stabilization via smooth feedback. In all scenarios, we present numerical examples to illustrate our results. To the best of our knowledge, our results provide the first theoretical link between the minimum-seeking approach introduced in [8], and the setting of synergistic hybrid control [20, 21].
The remainder of this manuscript is organized as follows. In section 2, we present the preliminaries. Section 3 presents the main problem formulation and main theoretical result. Section 4 specializes the theorem in section 3 to the problem of robust global stabilization on smooth manifolds and presents applications to two concrete problems: target-seeking with obstacle avoidance and rigid body attitude stabilization. Proofs are presented in section 5, followed by the conclusions and future work in section 6.
2 Preliminaries
2.1 Notation
We use $\langle x,y\rangle=x^{\top}y$, to denote the inner product between any two vectors $x,y\in\mathbb{R}^{n}$. Given a compact set $\mathcal{A}\subset\mathbb{R}^{n}$ and $x\in\mathbb{R}^{n}$, we use $|x|_{\mathcal{A}}:=\min_{\tilde{x}\in\mathcal{A}}\|x-\tilde{x}\|_{2}$. A set-valued mapping $M:\mathbb{R}^{p}\rightrightarrows\mathbb{R}^{n}$ is outer semicontinuous (OSC) at $z$ if for each sequence $\{z_{i},s_{i}\}\to(z,s)\in\mathbb{R}^{p}\times\mathbb{R}^{n}$ satisfying $s_{i}\in M(z_{i})$ for all $i\in\mathbb{Z}_{\geq 0}$, we have $s\in M(z)$. A mapping $M$ is locally bounded (LB) at $z$ if there exists an open neighborhood $N_{z}\subset\mathbb{R}^{p}$ of $z$ such that $M(N_{z})$ is bounded. The mapping $M$ is OSC and LB relative to a set $K\subset\mathbb{R}^{p}$ if $M$ is OSC for all $z\in K$ and $M(K):=\cup_{z\in K}M(x)$ is bounded. A function $\beta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ is of class ${K}\mathcal{L}$ if it is nondecreasing in its first argument, nonincreasing in its second argument, $\lim_{r\to 0^{+}}\beta(r,s)=0$ for each $s\in\mathbb{R}_{\geq 0}$, and $\lim_{s\to\infty}\beta(r,s)=0$ for each $r\in\mathbb{R}_{\geq 0}$. For two (or more) vectors $u,v\in\mathbb{R}^{n}$, we write $(u,v)=[u^{\top},v^{\top}]^{\top}$. If $x,y\in\mathbb{R}^{3}$, we use $[x]_{\times}$ to denote the the skew-symmetric matrix associated to $x$ and defined such that $[x]_{\times}y=x\times y$, where $\times$ indicates the cross product of vectors in $\mathbb{R}^{3}$. The notation $\otimes$ denotes the Kronecker product. Let $x=(x_{1},x_{2},\ldots,x_{n})\in\mathbb{R}^{n}$. Given a Lipschitz continuous function $f:\mathbb{R}^{n}\to\mathbb{R}^{m}$, we use $\partial_{x_{i}}f$ to denote the generalized Jacobian [25] of $f$ with respect to the variable $x_{i}$. A map $f$ is said to be of class $\mathcal{C}^{k}$ if it is $k$-times continuously differentiable with the $k$th-derivative being locally Lipschitz continuous. If $C\subset\mathbb{R}^{n}$, the notation $T_{x}C$ denotes the tangent cone [25] of $C$ at the point $x$. When $C$ is an embedded Euclidean submanifold, $T_{x}C$ denotes the tangent space at $x$, which is isomorphic to an affine linear subspace of $\mathbb{R}^{n}$.
2.2 Hybrid Dynamical Systems
In this paper, our models are given by hybrid dynamical systems (HDS), as studied in [26]. Such systems are characterized by the following inclusions:
:$\mathcal{H}:~{}~{}~{}\begin{cases}~{}~{}x\in C,&\dot{x}\hphantom{{}^{+}}\in F(x)\\ ~{}~{}x\in D,&x^{+}\in G(x),\end{cases}$ (1a)
where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is called the flow map, $G:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is called the jump map, $C\subset\mathbb{R}^{n}$ is called the flow set, and $D\subset\mathbb{R}^{n}$ is called the jump set. We use $\mathcal{H}=(C,F,D,G)$ to denote the data of the HDS $\mathcal{H}$. Purely continuous-time systems can be modeled as (1) by simply taking $D=\emptyset$. Similarly, purely discrete-time systems can be modeled as (1) by taking $C=\emptyset$. In this paper, we work with well-posed HDS that satisfy the following standing assumption.
Assumption 1.
The sets $C,D$ are closed. The set-valued mapping $F$ is OSC, LB, and for each $x\in C$ the set $F(x)$ is convex and nonempty. The set-valued mapping $G$ is OSC, LB, and for each $x\in C$ the set $G(x)$ is nonempty. $\square$
Henceforth, all hybrid systems in this manuscript are automatically assumed to satisfy the hybrid basic conditions.
Solutions to (1) are parameterized by a continuous-time index $t\in\mathbb{R}_{\geq 0}$, which increases continuously during flows, and a discrete-time index $j\in\mathbb{Z}_{\geq 0}$, which increases by one during jumps. Therefore, solutions to (1) are defined on hybrid time domains (HTDs). A set $E\subset\mathbb{R}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is called a compact HTD if $E=\cup_{j=0}^{J-1}([t_{j},t_{j+1}],j)$ for some finite sequence of times $0=t_{0}\leq t_{1}\ldots\leq t_{J}$. The set $E$ is a HTD if for all $(T,J)\in E$, $E\cap([0,T]\times\{0,\ldots,J\})$ is a compact HTD. The following definition formalizes the notion of solution to HDS of the form (1).
Definition 1.
A hybrid arc $x$ is a function defined on a HTD. In particular, $x:\text{dom}(x)\to\mathbb{R}^{n}$ is such that $x(\cdot,j)$ is locally absolutely continuous for each $j$ such that the interval $I_{j}:=\{t:(t,j)\in\text{dom}(x)\}$ has a nonempty interior. A hybrid arc $x:\text{dom}(x)\to\mathbb{R}^{n}$ is a solution $x$ to the HDS (1) if $x(0,0)\in C\cup D$, and: 1) For all $j\in\mathbb{Z}_{\geq 0}$ such that $I_{j}$ has nonempty interior: $x(t,j)\in C$ for all $t\in I_{j}$, and $\dot{x}(t,j)\in F(x(t,j))$ for almost all $t\in I_{j}$; 2) For all $(t,j)\in\text{dom}(x)$ such that $(t,j+1)\in\text{dom}(x)$: $x(t,j)\in D$ and $x(t,j+1)\in G(x(t,j))$. A solution $x$ is maximal if it cannot be further extended. A solution $x$ is said to be complete if $\text{length}~{}\text{dom}(x)=\infty$. $\square$
We will use the following standard stability definition:
Definition 2.
A compact set $\mathcal{A}\subset\mathbb{R}^{n}$ is said to be uniformly globally asymptotically stable (UGAS) for the HDS (1) if there exists $\beta\in\mathcal{KL}$ such that for each solution $x$ to (1) starting in $C\cup D$, we have
$|x(t,j)|_{\mathcal{A}}$ $\leq\beta(|x(0,0)|_{\mathcal{A}},t+j),$ (2)
for all $(t,j)\in\text{dom}(x)$. $\square$
In some cases, we consider HDS that depend on a small tunable parameter $\varepsilon\in\mathbb{R}_{>0}$, given by
$\mathcal{H}_{\varepsilon}:\begin{cases}C&\dot{x}\hphantom{{}^{+}}\in F_{\varepsilon}(x)\\ D&x^{+}\in G_{\varepsilon}(x).\end{cases}$ (3a)
For system (3), we use the following stability notion:
Definition 3.
[…]
defn:SPUAS For the HDS (3), a compact set $\mathcal{A}\subset\mathbb{R}^{n}$ is said to be Semi-Globally Practically Asymptotically Stable (SGpAS) as $\varepsilon\rightarrow 0^{+}$ if there exists $\beta\in\mathcal{KL}$ such that for each compact set $K\subset C\cup D$ and for each $\nu>0$, there exists $\varepsilon^{*}>0$ such that for all $\varepsilon\in(0,\varepsilon^{*}]$ and for all solutions of (3) with $x(0,0)\in K$, the following inequality holds:
$|x(t,j)|_{\mathcal{A}}\leq\beta(|x(0,0)|_{\mathcal{A}},t+j)+\nu,$ (4)
for all $(t,j)\in\text{dom}(x)$. $\square$
If there are several parameters that need to be tuned sequentially, we use the notation $(\varepsilon_{1},\varepsilon_{2},\dots,\varepsilon_{\ell})\rightarrow 0^{+}$ to encode the sequence of parameter tuning, i.e. $\varepsilon_{1}$ is tuned, then $\varepsilon_{2}$ after fixing $\varepsilon_{1}$, and so forth.
Finally, when $C\cup D$ is a compact set, SGpAS reduces to Uniform Global Practical Asymptotic Stability (UGpAS).
3 Strong $\nabla V$-Stabilizability of Control-Affine HDS
Consider the open-loop HDS:
$\begin{cases}\mathcal{X}_{C}&\dot{x}\hphantom{{}^{+}}=f(x,\theta,u)\\ \mathcal{X}_{D}&x^{+}\in G_{\mathcal{X}}(x)\end{cases}$ (5a)
where $x\in\mathbb{R}^{n}$, $u:=(u_{1},u_{2},\dots,u_{r})\in\mathbb{R}^{r}$ are the control inputs, $\mathcal{X}_{C}\subset\mathbb{R}^{n}$ is the flow set, $\mathcal{X}_{D}\subset\mathbb{R}^{n}$ is the jump set, $G_{\mathcal{X}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is the jump map, and $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{r}$ is the flow map, which has the following form:
$f(x,\theta,u)$ $=f_{0}(x,\theta)+\sum_{i=1}^{r}f_{i}(x,\theta)u_{i}.$ (5b)
In (5b), $\theta\in\Theta\subset\mathbb{R}^{m}$ represents a vector of potentially time-varying unknown parameters that models the unknown control directions. The class of systems (5) generalizes the class of control-affine systems commonly studied in the literature on continuous-time systems modeled as ODEs by allowing for jumps in the solutions. We also consider the class of systems (5b) as an open-loop HDS in the sense of [27]. In this context, the flow set, jump set, and discrete feedback law have been designed, and it remains to design the continuous feedback law.
The following regularity condition on (5) is needed to guarantee existence of maximal solutions and to rule out pathological behaviors, such as Zeno solutions.
Assumption 2.
For all $i\in\{1,2,\dots,r\}$ and all $(x,\theta)\in\mathcal{X}_{C}\times\Theta$, the following holds
  1. 1.
    $f_{0}$ and $f_{i}$ are $\mathcal{C}^{0}$, and $f_{i}(\cdot,\theta)$ is $\mathcal{C}^{1}$;
  2. 2.
    $f_{0}(x,\theta)\in T_{x}\mathcal{X}_{C}$, and $f_{i}(x,\theta)\in T_{x}\mathcal{X}_{C}$;
  3. 3.
    $G(\mathcal{X}_{D})\subset\mathcal{X}_{C}$, and $G(\mathcal{X}_{D})\cap\mathcal{X}_{D}=\emptyset$.
The vector of unknown parameters $\theta$ may be constant or time-varying. In the latter case, we require that $\theta$ is generated by an exogenous HDS of the form
$\begin{cases}\theta\in\Theta_{C}&\dot{\theta}\hphantom{{}^{+}}\in F_{e}(\theta)\\ \theta\in\Theta_{D}&\theta^{+}\in G_{e}(\theta)\end{cases}$ (6)
where $\Theta_{C},\Theta_{D}\subset\Theta$. We remark that the case of static uncertainty $\theta=\Theta$ is trivially included in (6) by taking $\Theta_{C}=\{\Theta\}$, $\Theta_{D}=\emptyset$, and $F_{e}(\theta)=\{0\}$. As usual, we require that the HDS (6) satisfies the hybrid basic conditions of Assumption 1. In addition, to guarantee existence of solutions to (6) from any initial condition in $\Theta_{C}\cup\Theta_{D}$, we impose the following regularity assumption.
Assumption 3.
$G_{e}(\Theta_{D})\subset\Theta_{C}$, and $G_{e}(\Theta_{D})\cap\Theta_{D}=\emptyset$.
Under Assumption 3, the HDS defined by (6) is sufficiently general to model a variety of complex behaviors. For example, the HDS (3) can model various classes of switching signals [26, Section 2.4], including those that have dwell-time bounds, which can be generated using an auxiliary hybrid automaton. While item (a) in Assumption 3 rules out consecutive jumps of $\theta$, this assumption can be relaxed to consider systems that satisfy average dwell-time bounds. However, to simplify our presentation, and since they rarely occur in practice, we rule out consecutive jumps triggered by the control direction.
Remark 1.
While $\theta$ is treated in this paper as an uncertain parameter, the regularity properties on the dynamics of $\theta$ are necessary so as to exclude pathological behaviors, e.g. Zeno solutions, purely discrete solutions, etc. In particular, by [28, Proposition 2.34], Assumption 3 guarantees that, due to the compactness of $\Theta$, there exists $t_{\circ}>0$ such that any solution $\theta$ to (6) satisfies $\overline{t}_{j}-\underline{t}_{j}\geq t_{\circ}>0$, where
$\overline{t}_{j}$ $:=\sup\{t\in\mathbb{R}_{\geq 0}~{}|~{}(t,j)\in\text{dom}(\theta)\},$
$\underline{t}_{j}$ $:=\,\inf\,\{t\in\mathbb{R}_{\geq 0}~{}|~{}(t,j)\in\text{dom}(\theta)\}.$
By interconnecting (5) and (6), we obtain the HDS
:$\mathcal{H}:\begin{cases}~{}~{}\xi\in C,&\dot{\xi}\hphantom{{}^{+}}\in F(\xi,u)\\ ~{}~{}\xi\in D,&\xi^{+}\in G(\xi)\end{cases}$ (7a)
where $\xi=(x,\theta)$, the sets $C$ and $D$ are given by
$C$ $=\mathcal{X}_{C}\times\Theta_{C},$ (7b)
$D$ $=(\mathcal{X}_{C}\times\Theta_{D})\cup(\mathcal{X}_{D}\times\Theta_{C})\cup(\mathcal{X}_{D}\times\Theta_{D}),$ (7c)
the flow map $F$ is given by
$F(\xi,u)$ =$=\{f(x,\theta,u)\}\times F_{e}(\theta),$ (7d)
and the jump map $G$ is given by
$G(\xi)=\begin{cases}G_{\mathcal{X}}(x)\times G_{e}(\theta)&(x,\theta)\in\mathcal{X}_{D}\times\Theta_{D}\\ \{x\}\times G_{e}(\theta)&(x,\theta)\in\mathcal{X}_{C}\times\Theta_{D}\\ G_{\mathcal{X}}(x)\times\{\theta\}&(x,\theta)\in\mathcal{X}_{D}\times\Theta_{C}\end{cases}$ (7e)
In words, the jump map (7e) allows jumps in the system whenever $x$, or $\theta$, are in their respective jump sets $\mathcal{X}_{D}$ and $\Theta_{D}$. Note that, by construction, $\mathcal{H}$ satisfies Assumption 1 whenever Assumptions 2 and 3 hold. In addition, the following Lemma is straightforward to verify.
Lemma 1.
Suppose that Assumptions 2 and 3 hold. Then, $G(D)\subset C$, and $G(D)\cap D=\emptyset$.
In the sequel, to simplify some expressions, we let $\mathcal{X}=\mathcal{X}_{C}\cup\mathcal{X}_{D}$, and we introduce the shorthand notation
$\dot{V}(\xi)$ $:=\langle\nabla V(x),\bar{f}(\xi)\rangle,$ (8a)
$\Delta V(\xi)$ $:=\max_{g\in G(\xi)}V(\pi_{x}(g))-V(\pi_{x}(\xi)),$ (8b)
where $\pi_{x}:\mathbb{R}^{n}\times\mathbb{R}^{m}\ni\xi=(x,\theta)\mapsto x\in\mathbb{R}^{n}$ is the canonical projection onto the first state, $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$ is a suitable $\mathcal{C}^{1}$ function to be characterized, and the map $\bar{f}$ is given by
$\bar{f}(\xi)$ $:=f_{0}(x,\theta)-\gamma\sum_{i=1}^{r}\langle\nabla V(x),f_{i}(x,\theta)\rangle f_{i}(x,\theta),$ (9)
where $\gamma>0$ is a constant. In what follows, $\mathcal{A}\subset\mathcal{X}$ is a compact subset. We now introduce the following definition, which makes use of the notation defined in (8).
Definition 3.1.
A $\mathcal{C}^{1}$ function $V:\mathbb{R}^{n}\rightarrow\mathbb{R}_{\geq 0}$ is said to be a Strong Control Lyapunov Function (SCLF) candidate with respect to $\mathcal{A}$ for $\mathcal{H}$ if there exists $\gamma>0$, and $\mathcal{K}_{\infty}$-functions $\alpha_{1}$, $\alpha_{2}$ such that
$\alpha_{1}(|x|_{\mathcal{A}})\leq$ $\,V(x)\leq\alpha_{2}(|x|_{\mathcal{A}}),$ $\forall x$ $\in\mathcal{X},$ (10)
$\dot{V}(\xi)\leq 0,$ $\forall\xi$ $\in C,$ (11)
$\Delta$ $V(\xi)\leq 0,$ $\forall\xi$ $\in D.$ (12)
Definition 3.1 is inspired by the notion of “strong $L_{g}V$ stabilizability” which was introduced in [8, 7]. Indeed, as discussed in [8, Sec. 3.1], the second condition in the Definition 3.1 can be written in a more explicit form as
$\langle\nabla V(x),f_{0}(x,\theta)\rangle-\gamma\sum_{i=1}^{r}\langle\nabla V(x),f_{i}(x,\theta)\rangle^{2}\leq 0,$ (13)
for all $(x,\theta)\in C$. If, in addition, for all $(x,\theta)\in C$ such that $|x|_{\mathcal{A}}=\epsilon>0$, the function $V$ satisfies the “strong small control property”:
$\lim_{\epsilon\rightarrow 0}\max_{\langle\nabla V(x),f_{0}(x,\theta)\rangle>0}\frac{\langle\nabla V(x),f_{0}(x,\theta)\rangle}{\sum_{i=1}^{r}\langle\nabla V(x),f_{i}(x,\theta)\rangle^{2}}<+\infty,$ (14)
then it is possible, see e.g. [7], to construct another SCLF candidate for which the non-increase inequality above is strengthened to a strict decrease:
$\langle\nabla V(x),f_{0}(x,\theta)\rangle-\gamma\sum_{i=1}^{r}\langle\nabla V(x),f_{i}(x,\theta)\rangle^{2}\leq-\alpha(|x|_{\mathcal{A}}),$
for some positive definite function $\alpha$. However, the new function will involve the expressions $\langle\nabla V(x),f_{i}(x,\theta)\rangle$, which necessitates the measurement of the unknown parameter $\theta$. Thus, we do not insist upon the strong small control property for $V$ so as to allow for some generality on how the unknown parameter $\theta$ affects the dynamics. Instead, if $V$ is an SCLF candidate with respect to $\mathcal{A}$ for $\mathcal{H}$, we associate to $V$ the HDS
$\mathcal{H}_{V}:\begin{cases}~{}~{}\xi\in C,&\dot{\xi}\hphantom{{}^{+}}\in\bar{F}(\xi)\\ ~{}~{}\xi\in D,&\xi^{+}\in G(\xi)\end{cases}$ (15)
wherein the flow map $\bar{F}$ is defined by
$\bar{F}(\xi)$ :=$:=\{\bar{f}(\xi)\}\times F_{e}(\theta),$
the map $\bar{f}$ is defined in (9), and the remaining data of $\mathcal{H}_{V}$ coincide with the data of $\mathcal{H}$. By construction, the compact subset $\mathcal{A}$ is Lyapunov stable for $\mathcal{H}_{V}$ [28, Theorem 3.18]. We then introduce the following definition.
Definition 3.2.
The HDS $\mathcal{H}$ is said to be strongly $\nabla V$-stabilizable if there exists an SCLF candidate $V$ with respect to $\mathcal{A}$ for $\mathcal{H}$ such that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$ is UGAS for $\mathcal{H}_{V}$. When $\mathcal{H}$ is strongly $\nabla V$-stabilizable, $V$ is said to be a Strong Control Lyapunov Function (SCLF) with respect to $\mathcal{A}$  for $\mathcal{H}$.
Remark 3.3.
In section 4, the state $x$ will be decomposed into $x=(p,q)$, where $p$ belongs to a smooth manifold $\mathcal{M}$ and $q$ belongs to a finite discrete set $\mathcal{Q}$. The state $q$ will play the role of a logic mode. In that case, the function $V$ will be given by $V(x)=V_{q}(p)$, where $\{V_{q}\}_{q\in\mathcal{Q}}$ is a family of smooth functions on $\mathcal{M}$ indexed by $q$. The main motivation in this case is the fact that, depending on the topology of $\mathcal{M}$, no single $\mathcal{C}^{1}$ function on $\mathcal{M}$ may satisfy the conditions of Definition 3.2, whereas a family of functions $\{V_{q}\}_{q\in\mathcal{Q}}$, indexed by the discrete state $q$, may.
Remark 3.4.
Establishing that $\mathcal{H}$ is strongly $\nabla V$-stabilizable is application dependent. The simplest case corresponds to the existence of a positive definite function $\alpha$ such that
$\dot{V}(\xi)$ $\leq-\alpha(|\pi_{x}(\xi)|_{\mathcal{A}}),$ $\forall\xi$ $\in C,$ (16a)
$\Delta V(\xi)$ $\leq-\alpha(|\pi_{x}(\xi)|_{\mathcal{A}}),$ $\forall\xi$ $\in D.$ (16b)
However, the inequalities (16) place significant restrictions on the unknown parameter $\theta$. In lieu of (16), we only require that $V$ is a weak Lyapunov function [28] for $\mathcal{H}_{V}$; i.e. that $V$ may be used in conjunction with the properties of the solutions of $\mathcal{H}_{V}$ to certify that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$ is UGAS for $\mathcal{H}_{V}$. This allows to model the practical situation wherein temporary loss of control inhibits the ability of feedback to induce strict decrease of $V$ during flows and/or jumps. We expand on this situation in section 4. $\square$
Remark 2.
The HDS $\mathcal{H}_{V}$ is obtained by closing the loop for the HDS $\mathcal{H}$ with the feedback law
$u_{i}(\xi)=-\gamma\langle\nabla V(x),f_{i}(\xi)\rangle,$ (17)
However, it is clear that the feedback law (17) necessitates real-time measurement of the state $x$ and the time-varying parameter $\theta$ in addition to complete knowledge of $V$ and $f_{i}$. Thus, the feedback law (17) cannot be implemented if the parameter $\theta$ is unknown, which is the situation of interest in this manuscript.
Assuming that $\mathcal{H}$ is strongly $\nabla V$-stabilizable, we propose a hybrid model-free feedback law that only requires real-time measurement or evaluations of $V(x)$. As shown in the stability analysis, under the proposed feedback law, the closed-loop system will behave approximately as system $\mathcal{H}_{V}$. This approximation property will be made precise below.
Specifically, we propose the feedback law
$u^{\varepsilon}_{i}(V(x),\eta)=\varepsilon^{-1}\sqrt{\frac{4\pi\gamma}{T_{i}\kappa}}\langle\exp(\kappa V(x)S)e_{1},\eta_{i}\rangle,$ (18a)
for $i\in\{1,2,\dots,r\}$, where $V$ is a SCLF with respect to $\mathcal{A}$ for $\mathcal{H}$, the constants $\kappa,\gamma,\varepsilon\in\mathbb{R}_{>0}$ are tuning parameters, $\eta_{i}\in\mathbb{S}^{1}$ for $i\in\{1,2,\dots,r\}$ is the state of an oscillator with the autonomous dynamics
$\dot{\eta}_{i}$ $=2\pi T_{i}^{-1}\varepsilon^{-2}S\eta_{i},$ $S$ $=\begin{pmatrix}0&1\\ -1&0\end{pmatrix},$ (18b)
and $\{T_{1},T_{2},\dots,T_{r}\}\subset\mathbb{Q}_{>0}$ is a collection constants such that $T_{i}=T_{j}$ if and only if $i=j$. For brevity, we will use
$\dot{\eta}$ $=\Lambda_{\varepsilon}(\eta),$ (18c)
to denote the collective (continuous-time) dynamics of the uncoupled oscillators $\eta=(\eta_{1},\eta_{2},\dots,\eta_{r})\in\mathbb{S}^{1}\times\mathbb{S}^{1}\times\cdots\times\mathbb{S}^{1}=\mathbb{T}^{r}$, where the map $\Lambda_{\varepsilon}$ on the right hand side of (18c) is defined consistently with (18b).
Closing the loop for $\mathcal{H}$ with the feedback law (18), we obtain the following HDS:
$\mathcal{H}_{cl}:\begin{cases}~{}~{}y\in C\times\mathbb{T}^{r},&\dot{y}\hphantom{{}^{+}}\in\hat{F}_{\varepsilon}(y)\\ ~{}~{}y\in D\times\mathbb{T}^{r},&y^{+}\in\hat{G}(y)\end{cases}$ (19a)
where, $y=(\xi,\eta)$, $\hat{G}(y)=G(\xi)\times\{\eta\}$, $\hat{F}_{\varepsilon}(y)=F_{\varepsilon}(\xi,\eta)\times F_{e}(\theta)$, and, for each $\varepsilon\in\mathbb{R}_{>0}$, the set-valued map $F_{\varepsilon}$ is
$F_{\varepsilon}(\xi,\eta)$ =$=\{f_{\varepsilon}(\xi,\eta))\}\times\{\Lambda_{\varepsilon}(\eta)\}\times F_{e}(\theta),$ (19b)
$f_{\varepsilon}(\xi,\eta)$ $=f_{0}(x,\theta)+\sum_{i=1}^{r}f_{i}(x,\theta)u^{\varepsilon}_{i}(V(x),\eta).$ (19c)
With $\bar{\mathcal{A}}:=\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{T}^{r}$, we can now state our first main result. The proof is presented in Section 5.
Theorem 3.
Let $\mathcal{H}$ be strongly $\nabla V$-stabilizable. Then, $\bar{\mathcal{A}}$ is SGpAS for $\mathcal{H}_{cl}$ as $\varepsilon\rightarrow 0^{+}$.
Naturally, we also have the following corollary.
Corollary 4.
If $\mathcal{X}$ is compact, then $\bar{\mathcal{A}}$ is UGpAS as $\varepsilon\rightarrow 0^{+}$ for $\mathcal{H}_{cl}$.
The proof of Theorem 3, presented in section 5, relies on the Lie-Bracket averaging theorems for well-posed hybrid dynamical systems with high amplitude high frequency oscillatory maps, which were recently introduced in [1]. We remark that, in contrast to the results in [1], which consider time-varying HDS, in the present paper we modeled the complete dynamics as time-invariant systems by using the dynamic oscillators (18c), which evolve in the compact set $\mathbb{T}^{r}$. In this way, Theorem 3 establishes stability of the compact subset $\bar{\mathcal{A}}$ for the HDS $\mathcal{H}_{cl}$, which includes $\mathbb{T}^{r}$. As a result, and given that the HDS satisfies Assumption 1, the stability property of Theorem 3 is robust in the sense of [26, Definition 7.18], i.e. it automatically inherits the robustness corollaries as formulated in [26, Chapters 7 & 9]. In particular, the following corollary is a direct consequence of this property.
Corollary 5.
Consider the perturbed closed-loop HDS $\mathcal{H}^{d}_{cl}$ (19), defined by
$\mathcal{H}^{d^{*}}_{cl}:\begin{cases}~{}~{}y+d_{1}\in C\times\mathbb{T}^{r},&\dot{y}\hphantom{{}^{+}}\in\hat{F}_{\varepsilon}(y+d_{2})+d_{3},\\ ~{}~{}y+d_{4}\in D\times\mathbb{T}^{r},&y^{+}\in\hat{G}(y+d_{5})+d_{6},\end{cases}$
where the signals $d_{i}:\text{dom}(y)\rightarrow(C\cup D)\times\mathbb{T}^{r}$ are any measurable functions satisfying
$\sup_{(t,j)\in\text{dom}(y)}|d_{i}(t,j)|\leq d^{*},$
for all $i\in\{1,2,\dots,6\}$ and some $d^{*}\in\mathbb{R}_{>0}$. Then, the perturbed HDS $\mathcal{H}^{d}_{cl}$ is SGpAS as $(d^{*},\varepsilon)\rightarrow 0^{+}$ and, if $\mathcal{X}$ is compact, then $\mathcal{H}^{d}_{cl}$ is UGpAS as $(d^{*},\varepsilon)\rightarrow 0^{+}$.
We note that the use of dynamic oscillators to capture oscillatory behaviors in hybrid seeking systems is not new, and it was also considered for the study of hybrid extremum seeking controller via first-order averaging [29, 30]. However, Theorem 3 is the first result that uses time-invariant oscillators for the analysis of Lie-bracket-based averaging algorithms.
4 Minimum Seeking for Synergistic Potential Functions on Smooth Manifolds
In this section, we present different applications of Theorem 3 in the context of robust global stabilization of a point $p^{\star}$ on a smooth manifold that is not globally contractible, and under unknown switching control gains. Henceforth, we assume that $\mathcal{M}$ is a smooth closed manifold properly embedded within, and equipped with the Riemannian metric of, an ambient Euclidean space $\mathbb{R}^{n_{p}}$.
Let $p\in\mathcal{M}$ denote the state of a plant whose dynamics is governed by the following control-affine system
$\dot{p}$ $=\sum_{i=1}^{r}b_{i}(p)\theta_{i}u_{i},$ (20)
where $u=(u_{1},u_{2},\dots,u_{r})$ are the control inputs. The vector $(\theta_{1},\theta_{2},\dots,\theta_{r})$ corresponds to unknown control gains that are allowed to dynamically switch between finitely many values, i.e. $(\theta_{1},\theta_{2},\dots,\theta_{r})\in\mathcal{E}\subset\mathbb{R}^{r}$ and $|\mathcal{E}|<+\infty$. For example, without loss of generality, we may consider the normalized directions =$\mathcal{E}=\{+1,0,-1\}^{r}$.
Remark 4.1.
For the purpose of illustration, in (20) we have specialized the form of uncertainty in the control directions to be linear. Nevertheless, we emphasize that general, potentially nonlinear, dependence on $\theta$ is allowed as long as Assumptions 2 and 3 hold.
To satisfy Assumption 3, we impose a dwell time condition on the rate of switching of $\theta$. We also require that each control gain $\theta_{i}$ does not vanish for an indefinite duration of time, since otherwise uniform (practical) asymptotic stability properties will be precluded. The two requirements are satisfied if $(\theta_{1},\theta_{2},\dots,\theta_{r})$ is governed by the HDS (6), where $\theta=(\theta_{1},\theta_{2},\dots,\theta_{r},\theta_{r+1},\theta_{r+2})\in\Theta_{C}\cup\Theta_{D}$, and the data of the system is given by
$\Theta_{C}$ :=×$:=\mathcal{E}\times[0,1]\times[0,T_{\circ}],$ (21a)
$\Theta_{D}$ :=×$:=\mathcal{E}\times\{1\}\times[0,T_{\circ}],$ (21b)
$F_{e}(\theta)$ :=$:=\{0\}\times\left[0,\chi_{1}\right]\times(\left[0,\chi_{2}\right]-\mathbb{I}_{\mathcal{E}_{b}}(\theta)),$ (21c)
$\mathcal{E}_{b}$ :=$:=\{\theta~{}|~{}\exists i\in\{1,2,\dots,r\}\text{ s.t. }\theta_{i}=0\},$ (21d)
$\tilde{G}_{e}(\theta)$ :=\$:=\mathcal{E}\backslash\{(\theta_{1},\theta_{2},\dots,\theta_{r})\},$ (21e)
$G_{e}(\theta)$ $:=\tilde{G}_{e}(\theta)\times\{0\}\times\{\theta_{r+2}\},$ (21f)
for some $T_{\circ}\in\mathbb{R}_{>0}$, $\chi_{1}\in\mathbb{R}_{>0}$, $\chi_{2}\in(0,1)$, and where $\mathbb{I}_{\mathcal{E}_{b}}(\cdot)$ is the classical indicator function on the subset $\mathcal{E}_{b}$.
The construction (21) involves a hybrid automaton, similar to the one considered in [31, Prop. 1.1], with auxiliary state $\theta_{r+1}\in\mathbb{R}_{\geq 0}$, and a time-ratio monitor with auxiliary state $\theta_{r+2}\in\mathbb{R}_{\geq 0}$, similar to the one considered in [29, Lemma 7]. In particular, every time the condition $\theta_{r+1}=1$ is satisfied, $\theta_{r+1}$ is reset to zero, and the vector of directions $\theta$ is allowed to jump according to the rule $\theta^{+}\in\tilde{G}_{e}(\theta)$, while $\theta_{r+2}$ remains unchanged. The role of $\theta_{r+2}$ is to model a monitor that decreases continuously whenever the current direction $\theta$ has a component equal to zero, i.e., when $\theta\in\mathcal{E}_{b}$ and therefore at least one of the entries in (20) has zero controllability. In this case, the condition $\theta_{r+2}=0$ will eventually occur, forcing $\theta$ to jump out of the set $\mathcal{E}_{b}$ and into the set $\{-1,1\}^{r}$. Note that, unlike existing hybrid monitors for switching systems with unstable modes [29], the set $\{-1,1\}^{r}$ also contains vectors $\theta$ with negative signs, which could be destabilizing for system (20) under a feedback law designed for $\theta=\mathbf{1}_{r}$. In other words, in this paper, the time-ratio monitor prevents the system from spending too much time in modes with zero control directions but not in modes with potentially problematic negative control directions.
By construction, and by [31, Prop. 1.1], and [29, Lemma 7], each hybrid arc generated by the HDS defined by (6) with data (21) satisfies the following dwell-time and activation-time inequalities for any two times $t_{2}>t_{1}$ in its domain:
$N_{\sharp}(t_{1},t_{2})$ $\leq\chi_{1}(t_{2}-t_{1})+1,$ (22a)
$T_{\sharp}(t_{1},t_{2})$ $\leq\chi_{2}(t_{2}-t_{1})+T_{\circ},$ (22b)
where $N_{\sharp}(t_{1},t_{2})$ is the total number of jumps during the time interval $(t_{1},t_{2})$, and
$T_{\sharp}(t_{1},t_{2})=\int_{t_{1}}^{t_{2}}\mathbb{I}_{\mathcal{E}_{b}}(\theta(t,j))\,\text{d}t,$ (23)
is the total time that the unknown vector $\theta$ spends in the set $\mathcal{E}_{b}$ during the same interval [29].
Next, we impose the following standard regularity condition on the control vector fields $b_{i}$ of system (20). This condition essentially guarantees that, in the ideal setting when $\theta=\mathbf{1}_{r}$, system (20) can actually be controlled by having vector fields $b_{i}$ that span the tangent space of the manifold $\mathcal{M}$ everywhere. Such assumptions are standard in the literature related to model-free stabilization and optimization [32, 7]. Similar assumptions have also been used in the design of synergistic hybrid feedback laws that overcome topological obstructions to stability [17].
Assumption 4.
For all $i\in\{1,2,\dots,r\}$ and all $p\in\mathcal{M}$, $b_{i}$ is $\mathcal{C}^{1}$, $b_{i}(p)\in T_{p}\mathcal{M}$, and there exists a constant $\lambda>0$ such that for all all $v\in T_{p}\mathcal{M}$, it holds that $\sum_{i=1}^{r}\langle b_{i}(p),v\rangle^{2}\geq\lambda\langle v,v\rangle$. $\square$
Let $V$ be a $\mathcal{C}^{1}$ function, and consider the subset
$\text{Crit}(V):=\{p\in\mathcal{M}~{}|~{}\nabla V(p)\in(T_{p}\mathcal{M})^{\perp}\},$ (24)
where $(T_{p}\mathcal{M})^{\perp}\subset\mathbb{R}^{n_{p}}$ denotes the orthogonal complement of the tangent space $T_{p}\mathcal{M}$ with respect to the Riemmannian metric of the ambient Euclidean space $\mathbb{R}^{n_{p}}$. With this notation, we recall the following definition, adapted from [20].
Definition 4.2.
For $N\in\mathbb{N}_{\geq 1}$ and $\mathcal{Q}=\{1,2,\dots,N\}$, a family of $\mathcal{C}^{1}$ functions $\{V_{q}\}_{q\in\mathcal{Q}}$, $V_{q}:\mathcal{M}\rightarrow\mathbb{R}_{\geq 0}$, is said to be a $\delta$-gap synergistic family of potential functions with respect to $p^{\star}$ if, for all $q\in\mathcal{Q}$, the following holds:
  1. 1.
    $V_{q}$ is positive definite with respect to $p^{\star}$;
  2. 2.
    $\forall c\in\mathbb{R}_{\geq 0}$, $\{p\in\mathcal{M}~{}|~{}V_{q}(p)\leq c\}$ is compact;
  3. 3.
    There exists an open neighborhood $\mathcal{U}_{q}\subset\mathbb{R}^{n_{p}}$ of $p^{\star}$ such that $\mathcal{U}_{q}\cap\text{Crit}(V_{q})=p^{\star}$;
  4. 4.
    There exists $\delta\in\mathbb{R}_{>0}$ such that $\delta<\Delta^{\star}$, where $\Delta^{\star}$ is
    $\Delta^{\star}=\min_{\begin{subarray}{c}q\in\mathcal{Q}\\ p\in\text{Crit}(V_{q})\end{subarray}}V_{q}(p)-\max_{\tilde{q}\in\mathcal{Q}}V_{\tilde{q}}(p).$ (25)
The construction of synergistic families of potential functions for stability problems typically requires qualitative information on the underlying manifold $\mathcal{M}$ and the target point $p^{\star}\in\mathcal{M}$. We make the assumption that such functions are available to us, and later we show how to satisfy this assumption in different applications.
Assumption 5.
The family of functions $\{V_{q}\}_{q\in\mathcal{Q}}$, $\mathcal{Q}=\{1,2,\dots,N\}$, is a $\delta$-gap synergistic family of potential functions with respect to $p^{\star}$.
To exploit the existence of synergistic potential functions, we introduce a logic state $q\in\mathcal{Q}$ into our controller, and we let $x=(p,q)\in\mathcal{M}\times\mathcal{Q}$, and define
$\mathcal{X}_{C}$ :=$:=\{x\in\mathcal{M}\times\mathcal{Q}~{}|~{}\mu(x)\leq\delta\},$ (26a)
$\mathcal{X}_{D}$ :=$:=\{x\in\mathcal{M}\times\mathcal{Q}~{}|~{}\mu(x)\geq\delta\},$ (26b)
$G_{\mathcal{X}}(x)$ :=$:=\{p\}\times\big{\{}\arg\min_{\tilde{q}\in\mathcal{Q}}V(p,\tilde{q})\big{\}},$ (26c)
where the function $\mu:\mathcal{M}\times\mathcal{Q}\rightarrow\mathbb{R}_{\geq 0}$ is defined by
$\mu(x):=V_{q}(p)-\textstyle\min_{\tilde{q}\in\mathcal{Q}}V_{\tilde{q}}(p).$ (26d)
Also, let $f_{0}$ and $f_{i}$ for $i\in\{1,2,\dots,r\}$ be given by
$f_{0}(x,\theta)$ $=0,$ $f_{i}(x,\theta)$ $:=\theta_{i}\begin{pmatrix}b_{i}(p)\\ 0\end{pmatrix}.$ (26e)
where the second entry in $f_{i}$ models the continuous-time dynamics of the logic state $q$, which remains constant during flows, i.e., $\dot{q}=0$. Finally, define the function $V:\mathcal{M}\times\mathcal{Q}\rightarrow\mathbb{R}_{\geq 0}$ as follows:
$V(x)=V_{q}(p).$ (27)
Then, we have the following proposition that applies to any system of the form (20) that satisfies the previous assumptions. The proof is presented in Section 5.
Proposition 6.
Suppose that Assumptions 4 and 5 are satisfied, and $V$ is given by (27). Then, $\mathcal{H}$ is strongly $\nabla V$-stabilizable. $\square$
The proof of Proposition 6 can be found in section 5. As a consequence of Theorem 3 and Proposition 6, we are able to conclude that, under the feedback law (18), the subset $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{T}^{r}$ is SGpAS as $\varepsilon\rightarrow 0^{+}$ for the closed loop HDS $\mathcal{H}_{cl}$ defined by (6), (7), (19), (21), and (26).
Figure 1: Global stabilization of $p^{\star}=(0,1)$ on $\mathbb{S}^{1}$ under unknown persistently switching control directions. Left: trajectories of (28) obtained under the proposed model-free feedback law (red color), and the non-hybrid model-free feedback law [8]. Right: Evolution in time of $\theta_{1}$.
We now present several concrete examples of the results in this section along with numerical simulation results. Henceforth, we take =$\mathcal{E}=\{+1,0,-1\}^{r}$. Since the construction of synergistic families of potential functions is not the main contribution of our manuscript, we rely on existing results in the literature. In particular, the functions used in the upcoming subsections are the same as those introduced in [20]. Alternative constructions, e.g. the universal construction technique introduced in [17], may also be used.
4.1 Robust Global Stabilization on $\mathbb{S}^{1}$
As the first example, we take $\mathcal{M}=\mathbb{S}^{1}$, $r=1$, and we consider the control-affine system
$\dot{p}$ $=b_{1}(p)\theta_{1}u_{1},$ $b_{1}(p)$ $=Sp$ (28)
where $p\in\mathbb{S}^{1}$, $S$ is the matrix defined in (18), $\theta_{1}\in\mathcal{E}=\{+1,0,-1\}$ is the unknown control gain, and $u_{1}$ is the control input. The goal is to globally stabilize a point $\vartheta^{\star}\in\mathbb{S}^{1}$. This problem was solved in [24] under the assumption of having constant and known control directions $\theta_{1}:=1$.
To globally stabilize $\vartheta^{\star}\in\mathbb{S}^{1}$ under unknown control directions, we consider the synergistic family of potential functions $\{W_{1},W_{2}\}$ defined by
$W_{q}(p)$ $:=W\circ\Phi_{q}(p),$ $W(p)$ $:=1-\langle\vartheta^{\star},p\rangle,$ (29a)
where $q\in\mathcal{Q}=\{1,2\}$, and the maps $\Phi_{q}:\mathbb{S}^{1}\rightarrow\mathbb{S}^{1}$ are
$\Phi_{q}(p):=\exp\left((3/2-q)W(p)S\right)p.$ (29b)
As shown in [24, 20], the family of functions $\{W_{1},W_{2}\}$ is a $\delta$-gap synergistic family of potential functions function with respect to $\vartheta^{\star}$ for any $\delta\in(0,1)$. It follows from Proposition 6 that the function $V$ defined by $V(p,q)=W_{q}(p)$ is an SCLF with respect to $\mathcal{A}=\{p^{\star}\}\times\mathcal{Q}$ for $\mathcal{H}$, and therefore that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{S}^{1}$ is SGpAS for $\mathcal{H}_{cl}$. Due to the compactness of $\mathbb{S}^{1}$, we may invoke Corollary 4 to conclude that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{S}^{1}$ is UGpAS.
Numerical simulations results illustrating the performance of the proposed controller for this example are shown in Figure 1. To generate the results, we used $\gamma=\sqrt{1}$, $\kappa=4$, $\varepsilon=1/\sqrt{4\pi}\approx 0.28$. We also used $\delta=1/4$ for the synergistic family of potential functions $\{W_{1},W_{2}\}$. The target point is $p^{\star}=(0,1)$. The right plot shows the evolution in time of the control direction $\theta$, which vanishes during bounded (but persistent) periods of time. Finally, to emphasize the robustness of the proposed feedback law, we added a small adversarial perturbation that locally stabilizes (in the absence of switching) the problematic critical point $p^{\sharp}=(0,-1)$. As shown in the figure, the proposed control law is not affected by the perturbation whereas a vanilla model-free feedback law is effectively trapped by the adversarial perturbation in the vicinity of the critical point $p^{\sharp}$.
Figure 2: Target-seeking with obstacle avoidance under unknown control gains. Left: trajectories of the vehicle under the proposed hybrid model-free feedback law (shown in red), and under the vanilla synergistic hybrid feedback (shown in blue) [17]. Right: Evolution in time of the control gain $\theta_{1}$. The gain $\theta_{2}=1$ is taken to be constant.
4.2 Robust Target Seeking with Obstacle Avoidance
Next, we consider the problem of stabilizing a target position for a mobile vehicle moving in an obstructed planar domain. Let $\mathcal{O}\subset\mathbb{R}^{2}$ denote the obstacle and let $z=(z_{1},z_{2})\in\mathbb{R}^{2}\backslash\mathcal{O}$ denote the position of the vehicle. The goal of the vehicle is to reach a target position $z^{\star}=(z_{1}^{\star},z_{2}^{\star})\in\mathbb{R}^{2}\backslash\mathcal{O}$ while avoiding the obstacle $\mathcal{O}$. We assume that the motion of the vehicle is governed by the kinematic equations
$\dot{z}$ $=\sum_{i=1}^{2}e_{i}\,\theta_{i}\,u_{i},$ (30)
where $u=(u_{1},u_{2})\in\mathbb{R}^{2}$ are the control input, and $(\theta_{1},\theta_{2})\in\mathcal{E}=\{+1,0,-1\}^{r}$ are the unknown control gains. Since the obstacle $\mathcal{O}$ is bounded, there exists $z_{\mathcal{O}}\in\mathbb{R}^{2}$ and $d_{0}>0$ such that $\mathcal{O}\subset z_{\mathcal{O}}+d\,\mathbb{B}$ for all $d\geq d_{0}$. To guarantee feasibility, we impose the following assumption on the target point $z^{\star}$.
Assumption 6.
$\exists~{}d^{\star}>d_{0}$ such that $z^{\star}\in\mathbb{R}^{2}\backslash(z_{\mathcal{O}}+d^{\star}\,\mathbb{B})$.
To handle the obstacle, and following the ideas of [17], we introduce the map $\varphi:\mathbb{R}^{2}\backslash(z_{\mathcal{O}}+d_{0}\,\mathbb{B}))\rightarrow\mathbb{R}\times\mathbb{S}^{1}$ given by
$\varphi(z)$ $=(\log(|z-z_{\mathcal{O}}|-d^{\star}),(z-z_{\mathcal{O}})/|z-z_{\mathcal{O}}|),$ (31)
As shown in [17], $\varphi$ is a well-defined diffeomorphism. The pushforward [33] of the kinematics of the vehicle under the diffeomorphism $\varphi$ are given by
$\dot{p}$ $=\sum_{i=1}^{2}b_{i}(p)\theta_{i}u_{i},$ $b_{i}(p)$ $=\text{D}\varphi\circ\varphi^{-1}(p)e_{i},$ (32)
where $p=(\rho,\vartheta)\in\mathbb{R}\times\mathbb{S}^{1}$. Therefore, global stabilization of the target position $z^{\star}$ in $\mathbb{R}^{2}\backslash(p_{\mathcal{O}}+d_{0}\,\mathbb{B}))$ is equivalent to globally stabilizing the point $p^{\star}=(\rho^{\star},\vartheta^{\star})$ on the smooth manifold =×$\mathcal{M}=\mathbb{R}\times\mathbb{S}^{1}$. However, the topology of $\mathcal{M}$ prohibits global stabilization by continuous feedback since $\mathbb{S}^{1}$ is a compact boundary-less manifold. With that in mind, we introduce the family of functions
$V_{q}(p)$ $:=\frac{1}{2}(\rho-\rho^{\star})^{2}+\sqrt{(\text{e}^{\rho}-\text{e}^{\rho^{\star}})^{2}+1}-1+W_{q}(\vartheta),$
where $q\in\mathcal{Q}=\{1,2\}$, and $W_{q}$ are the functions defined in (29). As in the previous subsection, the family of functions $\{W_{q}\}_{q\in\mathcal{Q}}$ is a $\delta$-gap synergistic Lyapunov function for $\vartheta^{\star}$ on $\mathbb{S}^{1}$ for any $\delta\in(0,1)$, which follows by [20]. Therefore, it is straightforward to show that the family of functions $\{V_{1},V_{2}\}$ is a $\delta$-gap synergistic family of potential functions with respect to $p^{\star}$. Thus, from Proposition 6, the function $V$ defined by $V(p,q)=V_{q}(p)$ is an SCLF with respect to $\mathcal{A}=\{p^{\star}\}\times\mathcal{Q}$ for $\mathcal{H}$. By invoking Theorem 3, we conclude that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{T}^{2}$ is SGpAS for the original hybrid system with unknown and dynamic control directions $\mathcal{H}_{cl}$.
To demonstrate the performance of the proposed controller compared to existing synergistic hybrid feedback [17], we provide numerical simulations results. The results are shown in Figure 5. To generate the results, we used $\gamma=2$, $\delta=1/4$, $\kappa=4$, and $\varepsilon=1/\sqrt{6\pi}\approx 0.165$. The target position is $z^{\star}=(0,2)$, and an obstacle with radius $d=1$ is centered at the origin, i.e. $z_{\mathcal{O}}=(0,0)$. Due to the special structure of the control vector fields in this example, we are able to use a single oscillator $\eta_{1}$ with period $T_{1}=1$ and rely on a $\pi/2$ phase shift to guarantee non-resonance between the inputs $u_{1}$ and $u_{2}$, see e.g. [34]. Figure 5 clearly indicates that the vanilla synergistic feedback fails to reach the target. This is to be expected since vanilla synergistic feedback requires knowledge of the control gain sign. By contrast, the proposed algorithm is able to reach the target despite the persistent switching in the control gain sign and the temporary loss of control.
Figure 3: Global stabilization of $p^{\star}=(0,0,1)$ on $\mathbb{S}^{2}$ under unknown persistently switching control directions. Left: trajectory corresponding to the proposed hybrid model-free feedback law (18) (shown in red), and the non-hybrid, model-free feedback law [8, 35] (shown in blue). Right: Evolution in time of the control gains $(\theta_{1},\theta_{2},\theta_{3})$.
4.3 Robust Global Stabilization on $\mathbb{S}^{2}$
For the third example, we take $\mathcal{M}=\mathbb{S}^{2}$, $r=3$, and we consider the control-affine system
$\dot{p}$ $=\sum_{i=1}^{3}b_{i}(p)\theta_{i}u_{i},$ $b_{i}(p)$ $=e_{i}-\langle p,e_{i}\rangle p,$ (33)
where $p\in\mathbb{S}^{2}$, $(\theta_{1},\theta_{2},\theta_{3})\in\mathcal{E}=\{+1,0,-1\}^{3}$ are the unknown control gains, and $u=(u_{1},u_{2},u_{3})\in\mathbb{R}^{3}$ are the control inputs. The goal is to globally stabilize an arbitrary point $p^{\star}\in\mathbb{S}^{2}$. To that end, we introduce the synergistic family of potential functions
$V_{q}(p)$ $:=W\circ\Phi_{q}(p),$ $W(p)$ $:=1-\langle p,p^{\star}\rangle,$ (34)
where $q\in\{1,2\}$, the maps $\Phi_{q}:\mathbb{S}^{2}\rightarrow\mathbb{S}^{2}$ are defined by
$\Phi_{q}(p):=\text{exp}\left((3/2-q)W(p)[p^{\star}_{\perp}]_{\times}\right)p,$ (35)
Figure 4: Global stabilization of $p^{\star}=vec(I)$ on $\text{SO}(3)$ under unknown switching control directions (shown in the bottom plot). The red trajectory corresponds to the proposed model-free feedback law (18), whereas the blue trajectory corresponds to a non-hybrid model-free feedback law [8]. Both systems are subject to the same small bounded persistent disturbance, which disrupts the blue trajectory.
and $p^{\star}_{\perp}\in\mathbb{S}^{2}$ is such that $\langle p^{\star}_{\perp},p^{\star}\rangle=0$. As shown in [20], the family of functions $\{V_{1},V_{2}\}$ is a $\delta$-gap synergistic family of potential functions function with respect to $p^{\star}$ for any $\delta\in(0,1)$. It follows from Proposition 6 that the function $V$ defined by $V(p,q)=V_{q}(p)$ is an SCLF with respect to $\mathcal{A}=\{p^{\star}\}\times\mathcal{Q}$ for $\mathcal{H}$, and therefore that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{S}^{2}$ is SGpAS for $\mathcal{H}_{cl}$. By invoking Corollary 4, we conclude that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{T}^{3}$ is UGpAS.
Numerical simulations results are shown in Figure 3. To generate the results, we used $\gamma=\sqrt{1}$, $\kappa=4$, $T_{1}=3$, $T_{2}=2$, $T_{3}=1$, $\delta=1/5$, and $\varepsilon=1/\sqrt{8\pi}$. The target point is $p^{\star}=(0,0,1)$, and $p_{\perp}^{\star}=(0,1,0)$. Finally, to emphasize the robustness of the proposed feedback law compared to vanilla model-free stabilization, e.g. [32], we added a small adversarial perturbation that locally stabilizes the problematic critical point $p^{\sharp}=(0,0,-1)$ in the absence of switching. As shown in Figure 3, the proposed control law is not affected by the perturbation whereas a vanilla model-free feedback law is trapped by the adversarial perturbation in the vicinity of the critical point $p^{\sharp}$.
4.4 Robust Global Stabilization on $\text{SO}(3)$
Next, we consider the problem of globally stabilizing a desired attitude $R^{\star}\in\text{SO}(3)$ for a rigid body. The kinematics of the rigid body are given by
$\dot{R}$ $=\sum_{i=1}^{3}R\,\widehat{e}_{i}\,\theta_{i}\,u_{i},$ (36)
where $R\in\text{SO}(3)\subset\mathbb{R}^{3\times 3}$ is the rotation matrix representing the attitude of the rigid body, $u=(u_{1},u_{2},u_{3})\in\mathbb{R}^{3}$ are the control inputs, and $\theta_{i}$ are the unknown control gains. Since $\text{SO}(3)$ is a Lie Group, the problem of stabilizing any specific attitude $R^{\star}\in\text{SO}(3)$ is equivalent to stabilizing the identity element, i.e. $I\in\text{SO}(3)$. Therefore, without loss of generality, we only consider the case when $R^{\star}=I$. We remark that, herein, we consider $\text{SO}(3)$ as an embedded submanifold of $\mathbb{R}^{3\times 3}$ equipped with its Riemannian metric, i.e. the Frobenius norm.
Following [20], we introduce the family of functions
$\tilde{V}_{q}(R)$ $:=W\circ\Phi_{q}(R),$ $W(R)$ $:=\text{tr}(A(I-R)),$ (37)
where $q\in\mathcal{Q}=\{1,2\}$, $A$ is the matrix given by
$A=\frac{3}{\sum_{i=1}^{3}\langle\tilde{\omega},e_{i}\rangle}\sum_{i=1}^{3}\langle\tilde{\omega},e_{i}\rangle e_{i},$ (38)
$\tilde{\omega}=(11,12,13)$, the maps $\Phi_{q}:\text{SO}(3)\rightarrow\text{SO}(3)$ are
$\Phi_{q}(R):=\exp\left(\frac{(3-2q)}{4}W(R)[\omega]_{\times}\right)R,$ (39)
and $\omega=\tilde{\omega}/\lVert\tilde{\omega}\lVert\in\mathbb{S}^{2}$. As shown in [20], the family of functions $\{\tilde{V}_{1},\tilde{V}_{2}\}$ is a $\delta$-gap synergistic family of potential functions with respect to $I$ for any $\delta\in(0,1/2)$.
Next, let $\mathcal{M}$ be the Euclidean submanifold
=$\mathcal{M}=\{{vec}(R)~{}|~{}R\in\text{SO}(3)\}\subset\mathbb{R}^{9}.$ (40)
By defining $p={vec}(R)\in\mathcal{M}$, it follows that $p$ evolves according to the driftless control-affine system
$\dot{p}$ $=\sum_{i=1}^{3}b_{i}(p)\,\theta_{i}\,u_{i},$ $b_{i}(p)$ $=-(\widehat{e}_{i}\otimes I)p$ (41)
Since the map ${vec}:\text{SO}(3)\rightarrow\mathcal{M}$ is a diffeomorphism, the family of functions $\{\tilde{V}_{1},\tilde{V}_{2}\}$ can be pulled back (via the inverse of $vec$) to a $\delta$-gap synergistic family of potential functions with respect to $p^{\star}={vec}(I)$. More explicitly, the family of functions $\{V_{1},V_{2}\}$ defined by $V_{q}(p)=\tilde{V}_{q}\circ vec^{-1}(p)$ is a $\delta$-gap synergistic family of potential functions with respect to $p^{\star}={vec}(I)$. Therefore, it follows from Proposition 6 that the function $V$ defined by $V(p,q)=V_{q}(p)$ is an SCLF with respect to $\mathcal{A}=\{p^{\star}\}\times\mathcal{Q}$ for the HDS $\mathcal{H}$, and therefore that $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})\times\mathbb{T}^{3}$ is SGpAS for $\mathcal{H}_{cl}$.
We now provide numerical simulations. The results of the simulations are shown in Figure 4. To generate the results, we used $\gamma=\sqrt{1}$, $\kappa=4$, $T_{1}=1$, $T_{2}=2$, $T_{3}=3$, $\delta=1/5$, and $\varepsilon=1/\sqrt{12\pi}$. The target point is $p^{\star}=vec(I)=(1,0,0,0,1,0,0,0,1)$. To emphasize the robustness of the proposed feedback law compared to vanilla model-free stabilization, e.g. [32], we added a small adversarial perturbation that locally stabilizes the (bad) critical point $p^{\sharp}=(-1,0,0,0,1,0,0,0,-1)$ in the absence of switching. As shown in Figure 4, the proposed control law is not affected by the perturbation whereas the vanilla model-free feedback law is trapped by the adversarial perturbation in the vicinity of the critical point $p^{\sharp}$.
4.5 Robust Target Seeking with Obstacle Avoidance for a Nonholonomic Vehicle
Finally, we consider again the problem of robust global stabilization of a target position for a mobile vehicle moving in an obstructed planar domain. However, in contrast to the single integrator dynamics considered in the previous section, we now consider a nonholonomic vehicle model. Due to the nonholonomic nature of the vehicle, the control vector fields do not span the entire tangent space of the state space manifold, and therefore Proposition 6 is not applicable. Although we are still able to apply the model-free feedback law (18), the stability analysis is different from the examples considered hitherto. Thus, we treat this special case separately.
Let $\mathcal{O}\subset\mathbb{R}^{2}$ denote the obstacle and $z=(z_{1},z_{2})\in\mathbb{R}^{2}\backslash\mathcal{O}$ denote the position of the vehicle. Suppose that the motion of the vehicle is governed by the nonholonomic kinematic equations
$\dot{z}$ $=\theta_{1}u_{1}\psi,$ $\dot{\psi}$ $=u_{2}S\psi$ (42)
where $u=(u_{1},u_{2})\in\mathbb{R}^{2}$ are the control inputs, $\theta_{1}\in\mathcal{E}\subset\mathbb{R}$ is the unknown control gain, $S$ is the matrix defined in (18b), and $\psi\in\mathbb{S}^{1}$ is a unit vector that indicates the forward direction of motion, i.e. the orientation, of the vehicle. This kinematic model has also been considered in [36, 37, 16, 38, 39, 40, 41, 35]. The goal of the vehicle is to stabilize a known target position $z^{\star}=(z_{1}^{\star},z_{2}^{\star})\in\mathbb{R}^{2}\backslash\mathcal{O}$, irrespective of the orientation $\psi\in\mathbb{S}^{1}$, while avoiding the obstacle $\mathcal{O}$. We impose Assumption 6 on the position of the target point $z^{\star}$ relative to the obstacle.
Figure 5: Simulation results for the problem of target-stabilization with obstacle avoidance for the nonholonomic kinematics (43).
Under the diffeomorphism defined by (31), the pushforward of the nonholonomic kinematics of the vehicle is given by
$\dot{p}$ $=\theta_{1}u_{1}\text{D}\varphi\circ\varphi^{-1}(p)\psi,$ $\dot{\psi}$ $=u_{2}S\psi,$ (43)
where $p=(\rho,\vartheta)\in\mathbb{R}\times\mathbb{S}^{1}$. Therefore, the goal of globally stabilizing the subset $\{z^{\star}\}\times\mathbb{S}^{1}$ in $\big{(}\mathbb{R}^{2}\backslash(z_{\mathcal{O}}+d_{0}\,\mathbb{B}))\big{)}\times\mathbb{S}^{1}$ is equivalent to globally stabilizing the compact subset $\{p^{\star}\}\times\mathbb{S}^{1}$ on the smooth manifold $\varphi\big{(}\mathbb{R}^{2}\backslash(z_{\mathcal{O}}+d_{0}\,\mathbb{B}))\big{)}\times\mathbb{S}^{1}=\mathcal{M}\times\mathbb{S}^{1}$, where $p^{\star}=(\rho^{\star},\vartheta^{\star})=\varphi(z^{\star})$. In contrast to the single integrator kinematics considered in subsection 4.2, Proposition 6 is not applicable for a vehicle with the nonholonomic kinematics (43) since the control vector fields in (43) clearly do not span the entire tangent space of the manifold $\mathcal{M}$ and so Assumption 4 is violated. Nevertheless, we will show that the structure of the control-affine system (43) permits the use of a slightly different model-free feedback law that stabilizes $\mathcal{H}$. To that end, we introduce the family of functions
$V_{q}(\mu)$ $:=\frac{1}{2}(\rho-\rho^{\star})^{2}+\sqrt{(\text{e}^{\rho}-\text{e}^{\rho^{\star}})^{2}+1}-1+W_{q}(\vartheta),$
where $q\in\mathcal{Q}=\{1,2\}$, and $W_{q}$ are the functions defined in (29). As in subsection 4.2, the family of functions $\{V_{1},V_{2}\}$ is a $\delta$-gap synergistic family of potential functions with respect to $\mu^{\star}=(\rho^{\star},\vartheta^{\star})$ on ×$\mathbb{R}\times\mathbb{S}^{1}$. Let
$\mu:=(\rho,\vartheta)\in\mathbb{R}\times\mathbb{S}^{1},$
$p=(\mu,\psi)\in\mathbb{R}\times\mathbb{S}^{1}\times\mathbb{S}^{1}=\mathcal{M}$, $x=(p,q)\in\mathcal{M}\times\mathcal{Q}$, and define the function $V:\mathcal{M}\times\mathcal{Q}\rightarrow\mathbb{R}_{\geq 0}$ as follows:
$V(x)=V_{q}(\mu).$ (44)
Also, let $r=1$, $T_{1}=1$, $\eta=\eta_{1}\in\mathbb{S}^{1}$, and let the first control input $u_{1}$ be given by the feedback law (18), i.e. we take $u_{1}$ to be
$u^{\varepsilon}_{1}(V(x),\eta)=\varepsilon^{-1}\sqrt{\frac{4\pi\gamma}{\kappa}}\langle\exp(\kappa V(x)S)e_{1},\eta_{1}\rangle.$ (45a)
Next, let the second control input $u_{2}$ to be defined as:
$u^{\varepsilon}_{2}=2\pi\varepsilon^{-1}.$ (45b)
Finally, define the compact sets $\mathcal{A}=\{\mu^{\star}\}\times\mathbb{S}^{1}\times\mathcal{Q}$, and $\bar{\mathcal{A}}=\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$. Then, we have the following proposition.
Proposition 7.
$\bar{\mathcal{A}}\times\mathbb{S}^{1}$ is SGpAS as $\varepsilon\rightarrow 0^{+}$ for $\mathcal{H}_{cl}$ with $V$ given by (44).
The proof of Proposition 7, presented in Section 5, relies on the multi-timescale Lie-bracket averaging tools developed by the authors in [1]. Specifically, by taking the second control input to be $u_{2}=2\pi\varepsilon^{-1}$, we induce an intermediate timescale $\tau_{1}=\varepsilon^{-1}t$ between the slow time scale $t$ and the fast timescale $\tau_{2}=O(\varepsilon^{-2}t)$. Then, using the recursive averaging results in [1], we average over both timescales $\tau_{1}$ and $\tau_{2}$, which allows to overcome the instantaneous lack of control authority.
We conclude this section with a numerical simulation result illustrating the control of the non-holonomic vehicle (42). The resulting trajectories of the vehicle are shown in Figure 5. To generate the results, we used $\gamma=2$, $\delta=1/4$, $\kappa=4$, and $\varepsilon=1/\sqrt{6\pi}\approx 0.165$. The target position is $z^{\star}=(0,2)$, and we considered an obstacle with radius $d=1$, centered at the origin, i.e. $z_{\mathcal{O}}=(0,0)$. The system is simulated under an additive adversarial perturbation designed to trap the trajectories of the vehicle behind the obstacle. The red trajectory corresponds to the proposed hybrid model-free feedback law, whereas the blue trajectory corresponds to a non-hybrid, model-free feedback law [8, 35]. The arrows indicate the direction of motion for the corresponding trajectory. The figure on the bottom depicts the magnitude of the adversarial perturbation which effectively traps the vanilla model-free feedback behind the obstacle. The figure on the top right depicts the control gain $\theta_{1}$ as a function of time. The figure on the top left depicts the obstructed planar domain wherein the vehicle operates.
5 Proofs
In this section, we present the proofs of our main results.
5.1 Proof of Theorem 3
Let $\tau\in\mathbb{R}_{\geq 0}$ and consider the HDS ~$\widetilde{\mathcal{H}}_{cl}$ with state $(\xi,\eta,\tau)\in\mathbb{R}^{n}\times\mathbb{R}^{r}\times\mathbb{R}_{\geq 0}$ and dynamics:
$\begin{cases}C\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0},&\begin{pmatrix}\dot{\xi}\\ \dot{\eta}\\ \dot{\tau}\end{pmatrix}\in F_{\varepsilon}(\xi,\eta)\times\{\Lambda_{\varepsilon}(\eta)\}\times\{\varepsilon^{-2}\}\\ D\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0},&\begin{pmatrix}\xi^{+}\\ \eta^{+}\\ \tau^{+}\end{pmatrix}\in G(\xi)\times\{\eta\}\times\{\tau\}\end{cases}$
where $F_{\varepsilon}$ is given by (19b), $\Lambda_{\varepsilon}(\eta)$ is given by (18c), $G$ is given by (7e), and $C,D$ are given by (7b). System ~$\widetilde{\mathcal{H}}_{cl}$ is a trivial dynamic extension of the closed-loop HDS $\mathcal{H}_{cl}$. Therefore, given an initial condition, any solution of the HDS $\mathcal{H}_{cl}$ corresponds to some solution of the HDS ~$\widetilde{\mathcal{H}}_{cl}$. Let $\zeta:=(\zeta_{1},\zeta_{2},\dots,\zeta_{r})\in\mathbb{S}^{1}\times\mathbb{S}^{1}\times\cdots\mathbb{S}^{1}=\mathbb{T}^{r}$ and consider the HDS ^$\widehat{\mathcal{H}}_{cl}$ defined by
$\begin{cases}C\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0},&\begin{pmatrix}\dot{\xi}\\ \dot{\zeta}\\ \dot{\tau}\end{pmatrix}\in\tilde{F}_{\varepsilon}(\xi,\zeta,\tau)\times\{0\}\times\{\varepsilon^{-2}\}\\ D\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0},&\begin{pmatrix}\xi^{+}\\ \zeta^{+}\\ \tau^{+}\end{pmatrix}\in G(\xi)\times\{\zeta\}\times\{\tau\}\end{cases}$ (46)
where $\tilde{F}_{\varepsilon}$ is defined by
$\tilde{F}_{\varepsilon}(\xi,\zeta,\tau):=F_{\varepsilon}(\xi,\exp(\Omega\tau)\zeta),$
and the matrix $\Omega$ is a block diagonal matrix of size $2r\times 2r$ with $r$ diagonal blocks of size $2\times 2$ such that the $i$-th block is the matrix $2\pi T_{i}^{-1}S$, where $S$ is the matrix defined in (18). Let $(\xi,\eta,\tau):\text{dom}(\xi,\eta,\tau)\rightarrow\mathbb{R}^{n}\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0}$ be any solution of ~$\widetilde{\mathcal{H}}_{cl}$ and define the hybrid arc $(\xi,\zeta,\tau):\text{dom}(\xi,\eta,\tau)\rightarrow\mathbb{R}^{n}\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0}$ by
$(\xi,\zeta,\tau):=(\xi,\exp(-\Omega\tau)\eta,\tau).$ (47)
Direct computation shows that
$\dot{\zeta}$ $=\exp(-\Omega\tau)\dot{\eta}-\dot{\tau}\exp(-\Omega\tau)\Omega\eta=0.$ (48)
Since, for any $\tau\in\mathbb{R}_{\geq 0}$ and any $\eta$ in $\mathbb{T}^{r}$, $\exp(-\Omega\tau)\eta\in\mathbb{T}^{r}$, it follows that the hybrid arc $(\xi,\zeta,\tau)$ defined by (48) is a solution of ^$\widehat{\mathcal{H}}_{cl}$ with the same initial condition for $\xi$ and $\tau$, and the following initial condition for $\zeta$:
$\zeta(0,0)=\exp(-\Omega\tau(0,0))\eta(0,0).$ (49)
Hence, for any solution of the HDS ~$\widetilde{\mathcal{H}}_{cl}$ there exists some solution of the HDS ^$\widehat{\mathcal{H}}_{cl}$ such that (48) holds. In particular, the $\xi$-component of both solutions coincide. We will use this equivalence to establish that every solution of ~$\widetilde{\mathcal{H}}_{cl}$ satisfies suitable stability bounds by first establishing such bounds for every solution of the HDS ^$\widehat{\mathcal{H}}_{cl}$.
An explicit computation of $\tilde{F}_{\varepsilon}$ shows that
$\tilde{F}_{\varepsilon}(\xi,\zeta,\tau)=\{\tilde{f}_{\varepsilon}(\xi,\zeta,\tau)\}\times F_{e}(\theta),$ (50)
where $F_{e}$ is given by (6), and
$\tilde{f}_{\varepsilon}(\xi,\zeta,\tau)=f_{0}(x,\theta)+\sum_{i=1}^{r}f_{i}(x,\theta)\hat{u}_{i}(V(x),\zeta_{i},\tau),$ (51)
and the functions $\hat{u}_{i}$ are given by
$\hat{u}_{i}(V,\zeta_{i},\tau)$ $=\varepsilon^{-1}\sqrt{\frac{4\pi\gamma}{T_{i}\kappa}}\langle\exp(\kappa VS)e_{1},\exp(2\pi S\tau/T_{i})\zeta_{i}\rangle.$
Now, since the constants $\{T_{1},T_{2},\dots,T_{r}\}$ are rational, it follows that there exists $T\in\mathbb{R}_{>0}$ such that the map $\tilde{f}_{\varepsilon}$ is $T$-periodic in $\tau$. Moreover, by definition of the map $\tilde{f}_{\varepsilon}$, the HDS ^$\widehat{\mathcal{H}}_{cl}$ belongs to the class of well-posed HDS with highly oscillatory flow maps which was analyzed in [1]. By applying the Lie-bracket averaging results in [1], we obtain that the Hybrid Lie-bracket averaged system ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$ corresponding to the HDS ^$\widehat{\mathcal{H}}_{cl}$ is given by
^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}:\begin{cases}~{}C\times\mathbb{T}^{r},&(\dot{\xi},\dot{\zeta})\hphantom{{}^{+}}\in\bar{F}(\xi)\times\{0\}\\ ~{}D\times\mathbb{T}^{r},&(\xi,\zeta)^{+}\in G(\xi)\times\{\zeta\}\end{cases},$ (52)
wherein the flow map $\bar{F}$ is defined by
$\bar{F}(\xi)$ :=$:=\{\bar{f}(x,\theta)\}\times F_{e}(\theta),$
$\bar{f}(\xi)$ $:=f_{0}(x,\theta)-\gamma{\sum_{i=1}^{r}}\langle\nabla V(x),f_{i}(x,\theta)\rangle f_{i}(x,\theta),$
which is independent of $\zeta$. Thus, the HDS ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$ is nothing but a trivial dynamic extension of the HDS $\mathcal{H}_{V}$ obtained by adding $\zeta$ as a state with trivial flow and jump dynamics, i.e., $\dot{\zeta}=0$ and $\zeta^{+}=\zeta$. From the assumptions of the theorem, namely that $\mathcal{H}$ is strongly $\nabla V$-stabilizable, it follows that the subset $\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$ is UGAS for the HDS $\mathcal{H}_{V}$ defined in (15). It follows that the compact subset $\bar{\mathcal{A}}$ is UGAS for the HDS ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$ given by (52). From [1, Theorem 2], we obtain that there exists a class $\mathcal{KL}$ function $\beta$ such that for each compact subset $K\subset(C\cup D)\times\mathbb{T}^{r}$ and for each $\nu>0$, there exists $\varepsilon^{*}>0$ such that for all $\varepsilon\in(0,\varepsilon^{*}]$ and for all solutions of ^$\widehat{\mathcal{H}}_{cl}$ with $(\xi(0,0),\zeta(0,0))\in K$, the following inequality holds for all $(t,j)\in\text{dom}(x,\zeta,\tau)$:
$|(\xi(t,j),\zeta(t,j))|_{\bar{\mathcal{A}}}\leq\beta(|(\xi(0,0),\zeta(0,0))|_{\bar{\mathcal{A}}},t+j)+\nu.$
With $\varepsilon\in(0,\varepsilon^{*}]$, let $(\xi,\eta,\tau):\text{dom}(\xi,\eta,\tau)\rightarrow\mathbb{R}^{n}\times\mathbb{T}^{r}\times\mathbb{R}_{\geq 0}$ be any solution of the HDS ~$\widetilde{\mathcal{H}}_{cl}$ such that $(\xi(0,0),\eta(0,0))\in K$. By construction, it follows that $(\xi,\eta,\tau)$ satisfies
$(\xi,\eta,\tau)=(\xi,\exp(\Omega\tau)\zeta,\tau),$ (53)
for some solution $(\xi,\zeta,\tau)$ of the HDS ^$\widehat{\mathcal{H}}_{cl}$ with $(\xi,\zeta)(0,0)\in K$. Therefore, we obtain that there exists a class $\mathcal{KL}$ function $\beta$ such that for each compact subset $K\subset(C\cup D)\times\mathbb{T}^{r}$ and for each $\nu>0$, there exists $\varepsilon^{*}>0$ such that for all $\varepsilon\in(0,\varepsilon^{*}]$ and for all solutions of ~$\widetilde{\mathcal{H}}_{cl}$ with $(\xi(0,0),\eta(0,0))\in K$, the following inequality holds for all $(t,j)\in\text{dom}(x,\eta,\tau)$:
$|(\xi(t,j),\eta(t,j))|_{\bar{\mathcal{A}}}\leq\beta(|(\xi(0,0),\eta(0,0))|_{\bar{\mathcal{A}}},t+j)+\nu.$
However, since ~$\widetilde{\mathcal{H}}_{cl}$ is nothing but a trivial dynamic extension of the closed loop system $\mathcal{H}_{cl}$, it follows that the $\mathcal{KL}$ bound above is also true for every solution of $\mathcal{H}_{cl}$ with $(\xi(0,0),\eta(0,0))\in K$. In particular, we may always take $K=K_{1}\times\mathbb{T}^{r}$ for some compact $K_{1}\subset(C\cup D)$ so that any initial condition $\eta(0,0)$ for the state of the oscillators is admissible. This concludes the proof.
5.2 Proof of Proposition 6
When $f_{0}$ and $f_{i}$ are as given in (26), direct computations show that $\bar{f}$, defined in (9), takes the form
$\bar{f}(\xi)=-\gamma\sum_{i=1}^{r}\theta_{i}^{2}\langle\nabla V_{q}(p),b_{i}(p)\rangle\begin{pmatrix}b_{i}(p)\\ 0\end{pmatrix}.$
As a result, if $\xi=(x,\theta)=((p,q),\theta)$, we have that
$\dot{V}(\xi)\leq-\gamma\sum_{i=1}^{r}\theta_{i}^{2}\langle\nabla V_{q}(p),b_{i}(p)\rangle^{2}\leq 0,$
for all $\xi\in C$. On the other hand, by construction we have that for all $\xi\in D$:
$\Delta V(\xi)\leq 0,$
Thus, $V$ is an SCLF candidate with respect to $\mathcal{A}$ for the HDS $\mathcal{H}$ defined by (5)-(7), (21), and (26).
Since $\mathcal{M}$ is equipped with the Riemannian metric of the ambient space and $b_{i}(p)\in T_{p}\mathcal{M}$, we have that
$\langle\nabla V_{q}(p),b_{i}(p)\rangle=\langle\nabla_{\mathcal{M}}V_{q}(p),b_{i}(p)\rangle,$
where $\nabla_{\mathcal{M}}V_{q}(p)$ is the unique orthogonal projection of $\nabla V_{q}(p)$ onto the tangent space $T_{p}\mathcal{M}$ with respect to the Riemannian metric of the ambient Euclidean space. Therefore, we obtain that
$\dot{V}(\xi)=-\gamma\sum_{i=1}^{r}\theta_{i}^{2}\langle\nabla_{\mathcal{M}}V_{q}(p),b_{i}(p)\rangle^{2}.$
On the other hand, by definition, if $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, then $\theta_{i}^{2}>0$, $\forall i\in\{1,2,\dots,r\}$. Since $\mathcal{E}$ is a discrete finite set, there exists a constant $\lambda_{1}\in\mathbb{R}_{>0}$ such that, for all $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we have $\theta_{i}^{2}>\lambda_{1}$ $\forall i\in\{1,2,\dots,r\}$.
Hence, for all $\xi\in C$ such that $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we have $\dot{V}(\xi)\leq-\gamma\lambda_{1}\sum_{i=1}^{r}\langle\nabla_{\mathcal{M}}V_{q}(p),b_{i}(p)\rangle^{2}$. From Assumption 4, there exists a constant $\lambda\in\mathbb{R}_{>0}$ such that, for all $p\in\mathcal{M}$ and all $v\in T_{p}\mathcal{M}$, we have that $\sum_{i=1}^{r}\langle v,b_{i}(p)\rangle^{2}\geq\lambda\langle v,v\rangle$. Hence, for all $\xi\in C$ such that $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we have
$\dot{V}(\xi)\leq-\gamma\lambda_{1}\lambda\langle\nabla_{\mathcal{M}}V_{q}(p),\nabla_{\mathcal{M}}V_{q}(p)\rangle.$
By construction of the family $\{V_{q}\}_{q\in\mathcal{Q}}$, we have that
$\nabla_{\mathcal{M}}V_{q}(p)=0,\iff p=p^{\star},$
for all $(p,q)\in\mathcal{X}_{C}\times\mathcal{Q}$. It follows that there exists a positive definite function $\rho$ such that
$\langle\nabla_{\mathcal{M}}V_{q}(p),\nabla_{\mathcal{M}}V_{q}(p)\rangle\geq\rho(|(p,q)|_{\mathcal{A}}),$
for all $(p,q)\in\mathcal{X}_{C}$, which implies that for all $\xi\in C$ such that $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we have that
$\dot{V}(\xi)\leq-\gamma\lambda_{1}\lambda\rho(|x|_{\mathcal{A}}).$
Following similar steps as in [42], it can be shown that there exists a continuously differentiable $\mathcal{K}_{\infty}$ function $\bar{\alpha}$ such that the function $\hat{V}=\bar{\alpha}\circ V$ satisfies
$\bar{\alpha}_{1}(|x|_{\mathcal{A}})\leq$ $\hat{V}(x)\leq\bar{\alpha}_{2}(|x|_{\mathcal{A}}),$ $\forall x$ $\in\mathcal{X},$
$\dot{\hat{V}}(\xi)\leq 0,$ $\forall\xi$ $\in C,$
$\Delta$ $\hat{V}(\xi)\leq 0,$ $\forall\xi$ $\in D.$
and, for all $\xi\in C$ such that $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we have
$\dot{\hat{V}}(\xi)\leq-\hat{V}(\xi).$
We now introduce the function $\tilde{V}(\xi)=\hat{V}(\xi)\text{e}^{\theta_{r+2}}$. Direct computations shows that, during flows, we have
$\dot{\tilde{V}}(\xi)$ $=(\dot{\hat{V}}(\xi)+\dot{\theta}_{r+2}\hat{V}(\xi))\text{e}^{\theta_{r+2}}$
From (21), we have that $\dot{\theta}_{r+2}\in[0,\chi_{2}]-\mathbb{I}_{\mathcal{E}_{b}}(\theta)$. Therefore, if $\theta\in\mathcal{E}_{b}$, we have that
$\dot{\theta}_{r+2}\leq-(1-\chi_{2})\implies\dot{\tilde{V}}(\xi)\leq-(1-\chi_{2})\tilde{V}(\xi).$
On the other hand, if $\theta\in(\Theta_{C}\cup\Theta_{D})\backslash\mathcal{E}_{b}$, we also have that
$\dot{\theta}_{r+2}\leq\chi_{2}\implies\dot{\tilde{V}}(\xi)\leq-(1-\chi_{2})\tilde{V}(\xi)$
From (21), we have that $(\theta_{r+2})^{+}=\theta_{r+2}$, for all $\xi\in D$. Thus, using the facts that $\theta_{r+2}\in[0,T_{\circ}]$ and $|x|_{\mathcal{A}}=|\xi|_{\tilde{\mathcal{A}}}$ for all $\xi\in C\cup D$, we obtain
$\bar{\alpha}_{1}(|\xi|_{\tilde{\mathcal{A}}})\leq$ $\tilde{V}(\xi)\leq\bar{\alpha}_{2}(|\xi|_{\tilde{\mathcal{A}}})\text{e}^{T_{\circ}},$ $\forall\xi$ $\in C\cup D,$ (54a)
$\dot{\tilde{V}}(\xi)\leq$ $-(1-\chi_{2})\tilde{V}(\xi),$ $\forall\xi$ $\in C,$ (54b)
$\Delta$ $\tilde{V}(\xi)\leq 0,$ $\forall\xi$ $\in D,$ (54c)
where $\tilde{\mathcal{A}}=\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$. Because the set $\tilde{\mathcal{A}}$ is Lyapunov stable for $\mathcal{H}_{V}$, which is entailed by the conditions (54) [43, Theorem 3.18], it follows that any maximal solution to $\mathcal{H}_{V}$ starting in $C\cup D$ is bounded. Thanks to Lemma 1, we obtain that, by invoking [43, Proposition 2.34], no nontrivial maximal solution to $\mathcal{H}_{V}$ starting in $C\cup D$ is Zeno and that, for every maximal solution $\xi$ to $\mathcal{H}_{V}$ starting in $C\cup D$, there exists $t_{\circ}>0$ such that $\overline{t}_{j}-\underline{t}_{j}\geq t_{\circ}>0$, for all $j\geq 1$, where $\overline{t}_{j}:=\sup\{t\in\mathbb{R}_{\geq 0}~{}|~{}(t,j)\in\text{dom}(\xi)\}$, and $\underline{t}_{j}:=\,\inf\,\{t\in\mathbb{R}_{\geq 0}~{}|~{}(t,j)\in\text{dom}(\xi)\}$. In other words, no solution of $\mathcal{H}_{V}$ is discrete. By invoking [43, Theorem 3.19], we obtain that the compact subset $\tilde{\mathcal{A}}$ is globally asymptotically stable in the sense of [43, Definition 3.1]. However, since $\tilde{\mathcal{A}}$ is compact and $\mathcal{H}_{V}$ is a well-posed HDS, it follows from [43, Theorem 3.22] that the global asymptotic stability of $\tilde{\mathcal{A}}$ is uniform, i.e. that $\tilde{\mathcal{A}}$ is UGAS for $\mathcal{H}_{V}$ in the sense of Definition 2 which concludes the proof.
5.3 Proof of Proposition 7
The closed-loop HDS $\mathcal{H}_{cl}$ is defined by (7) in conjunction with (6), (18), (19), (21), (26), (43), and (45). Let $\tau_{1},\tau_{2}\in\mathbb{R}_{\geq 0}$, $\tau=(\tau_{1},\tau_{2})\in\mathbb{R}_{\geq 0}^{2}=\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}$, and define the HDS ~$\widetilde{\mathcal{H}}_{cl}$ via the following data:
$\begin{cases}~{}\tilde{C}\times\mathbb{R}^{2}_{\geq 0},&\begin{pmatrix}\dot{\xi}\\ \dot{\eta}\\ \dot{\tau}\end{pmatrix}\in F_{\varepsilon}(\xi,\eta)\times\{\Lambda_{\varepsilon}(\eta)\}\times\left\{(\varepsilon^{-1},\varepsilon^{-2})\right\}\\ ~{}\tilde{D}\times\mathbb{R}_{\geq 0}^{2},&\begin{pmatrix}\xi^{+}\\ \eta^{+}\\ \tau^{+}\end{pmatrix}\in G(\xi)\times\{\eta\}\times\{\tau\}\end{cases}$
where $\tilde{C}=C\times\mathbb{T}^{1}$, $\tilde{D}=D\times\mathbb{T}^{1}$. It is clear that ~$\widetilde{\mathcal{H}}_{cl}$ is a trivial dynamic extension of $\mathcal{H}_{cl}$. Next, similar to the proof of Theorem 3, we introduce the HDS ^$\widehat{\mathcal{H}}_{cl}$ defined by
$\begin{cases}\tilde{C}\times\mathbb{R}^{2}_{\geq 0}&\begin{pmatrix}\dot{\mu}\\ \dot{\zeta}\\ \dot{\tau}\end{pmatrix}\in\widehat{F}_{\varepsilon}(\mu,\zeta,\tau)\times\{0\}\times\left\{(\varepsilon^{-1},\varepsilon^{-2})\right\}\\ \tilde{D}\times\mathbb{R}_{\geq 0}^{2}&\begin{pmatrix}\mu^{+}\\ \zeta^{+}\\ \tau^{+}\end{pmatrix}\in G(\mu)\times\{\zeta\}\times\{\tau\}\end{cases}$
where $\zeta\in\mathbb{S}^{1}$, $\mu=(p,\phi,q,\theta)\in\mathcal{M}\times\mathbb{S}^{1}\times\mathcal{Q}\times(\Theta_{C}\cup\Theta_{D})$, and the map $\widehat{F}_{\varepsilon}$ is given by
$\widehat{F}_{\varepsilon}(\mu,\zeta,\tau)=F_{\varepsilon}((p,\exp(2\pi\tau_{1}S)\phi,q,\theta),\exp(2\pi\tau_{2}S)\eta).$
An explicit computation shows that $\widehat{F}_{\varepsilon}(\mu,\zeta,\tau)=\{\widehat{f}_{\varepsilon}(\mu,\zeta,\tau)\}\times F_{e}(\theta)$, where the map $\widehat{f}_{\varepsilon}$ is defined by
$\widehat{f}_{\varepsilon}(\mu,\zeta,\tau)=\theta_{1}b_{1}(p,\phi,\tau_{1})\hat{u}_{1}^{\varepsilon}(V(x),\zeta,\tau_{2}),$ (55)
where the vector field $b_{1}$ is given by
$f_{1}(p,\phi,\tau_{1}):=\begin{pmatrix}\text{D}\varphi\circ\varphi^{-1}(p)\exp(2\pi\tau_{1}S)\phi\\ 0\\ 0\end{pmatrix},$ (56)
and the function $\hat{u}_{1}^{\varepsilon}$ is given by
$\hat{u}_{1}^{\varepsilon}(V,\zeta,\tau_{2})$ $=\varepsilon^{-1}\sqrt{\frac{2\gamma}{\kappa}}\langle\exp(\kappa VS)e_{1},\exp(2\pi S\tau_{2})\zeta\rangle.$
It is clear that $\widehat{f}_{\varepsilon}$ is $1$-periodic in $\tau_{1}$ and $\tau_{2}$. Moreover, the structure of the HDS ^$\widehat{\mathcal{H}}_{cl}$ is the same as the structure of the class of well-posed HDS with highly oscillatory flow maps considered in [1]. By applying the recursive Lie-bracket averaging formulas in [1], it can be shown that the Hybrid Lie-bracket averaged system ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$ corresponding to ^$\widehat{\mathcal{H}}_{cl}$ is the system defined by
$\begin{cases}(\mu,\zeta)\in\tilde{C},&(\dot{\mu},\dot{\zeta})\hphantom{{}^{+}}\in\bar{F}(\mu)\times\{0\}\\ (\mu,\zeta)\in\tilde{D},&(\mu,\zeta)^{+}\in G(\mu)\times\{\zeta\},\end{cases}$
where the map $\bar{F}(\mu,\zeta)=\{\bar{f}(\mu)\}\times F_{e}(\theta)$,
$\bar{f}(\mu)$ $=-\frac{\theta_{1}^{2}\gamma}{2}\sum_{i=1}^{2}\langle b_{i}(p),\nabla V_{q}(p)\rangle\begin{pmatrix}b_{i}(p)\\ 0\\ 0\end{pmatrix},$ (57)
and the maps $b_{i}$ are given by $b_{i}(p)=\text{D}\varphi\circ\varphi^{-1}(p)e_{i}$. It is clear that ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$ is a trivial dynamic extension of the HDS
$\mathcal{H}_{V}:\begin{cases}~{}\mu\in C,&~{}\dot{\mu}\hphantom{{}^{+}}\in\bar{F}(\mu)\\ ~{}\mu\in D,&~{}\mu^{+}\in G(\mu)\end{cases}$
Now, consider the function $V(\mu)=V_{q}(p)$ and observe that
$\dot{V}(\mu)=-\frac{\theta_{1}^{2}\gamma}{2}\sum_{i=1}^{2}\langle b_{i}(p),\nabla V_{q}(p)\rangle^{2}\leq 0,$ (58)
for all $\mu\in D$, and $\Delta V(\mu)\leq 0$, for all $\mu\in D$. In addition, $V$ is positive definite with respect to $\bar{\mathcal{A}}=\mathcal{A}\times(\Theta_{C}\cup\Theta_{D})$ and, for every $c\in\mathbb{R}_{\geq 0}$, the set $\{\mu\in C\cup D~{}|~{}V(\mu)\leq c\}$ is compact. Following similar steps to the proof of Proposition 6, it can be shown that $\bar{\mathcal{A}}$ UGAS for the HDS $\mathcal{H}_{V}$, which implies that $\bar{\mathcal{A}}\times\mathbb{S}^{1}$ is UGAS for the HDS ^$\widehat{\mathcal{H}}_{cl}^{\text{ave}}$. Then, following similar steps as in the proof of Theorem 3, we obtain that $\bar{\mathcal{A}}\times\mathbb{S}^{1}$ is SGpAS for $\mathcal{H}_{cl}$ as $\varepsilon\rightarrow 0^{+}$, which concludes the proof.
6 Conclusions
We studied the problem of robust global stabilization for a class of control-affine systems under dynamic uncertainty in the control directions and topological obstructions. By proposing a novel class of hybrid and highly-oscillatory feedback laws that seek the minimum of a family of synergistic Lyapunov functions, we provide a robust solution to this problem. Our method is particularly advantageous as it is model-free and can handle unknown and switching control gain signs, ensuring resilience in autonomous systems facing adversarial spoofing attacks. The practical relevance of our results is demonstrated through several concrete applications. Since our results allow to transfer stability properties from well-posed stable hybrid systems with known control directions to systems with unknown control directions, they open the door for the systematic development of new “model-free” controllers that can exploit the rich set of tools developed during the last two decades for hybrid systems. Numerical simulations validate the effectiveness and robustness of the proposed feedback law compared to existing techniques. Future research directions will focus on experimental validations of the proposed algorithms.